Consider the following data: 3, 10, 8, 4, 5, 1 Compute (1) the sample standard derivation; (2) the sample median: (3) the IQR: and (4) draw a boxplot Please round your answer to the first decimal point.
Answer #1

The sample can be re written as

1, 3, 4, 5, 8, 10

1) To calculate sample standard deviation we need to find mean first using

LX, 1+3+4 +5 +8+10 6 = 5.2

Now sample standard deviation calculated as

Σ(X-X)2 (5.167-1)2 + (5.167-3)2 + + (5.167-102 rl 6-1

2) .The median is the center value of o any distribution if the distribution is in ascending order hence here,

(5+4)/2= 4.5 is the median.

3) IQR = Inter quartile range , it is calculated by first calculating 5 pint summary, so

Q1=3, Q2=4.5, Q3= 8 now IQR= Q3-Q1= 8-3=5

4. The box plot

0595857565554535251 9 7